OpenStudy (anonymous):

3 - 5(x + 2) = 3(x + 1) + 6

OpenStudy (anonymous):

simplify,multiply, add, subtract then collect like terms

OpenStudy (anonymous):

0_o thnks?

OpenStudy (anonymous):

3 - 5(x + 2) = 3(x + 1) + 6 3-5x -10 = 3x+ 3 +6 -5x -7 = 3x +9 -8x=16 x= 16/-8 x=-2

OpenStudy (anonymous):

thnks bro

OpenStudy (anonymous):

ok yw